# OUJ-FTC-21 OCHA-PP-384

### v3

@@ 3.2 Hindmarsh-Rose model
$\frac{\partial^{2}H_{1}}{\partial X^{2}}+\frac{\partial^{2}H_{1}}{\partial Y^{2}}$ $=$ $1+2dX,$ (65)
@@
$\frac{\partial H_{1}}{\partial X}$ $=$ $+f(X,Y),$ (68)
@@
$\frac{\partial f(X,Y)}{\partial X}+\frac{\partial g(X,Y)}{\partial X}$ $=$ $1+2dX.$ (70)
@@
About $H_{2}$ increasing, $H_{2}=z,\ \dot{H}_{2}=r\{\,s(x_{R}-x)+z\,\},$ with $r\ll 1$, so that $z$ acts as a slow variable. The sign of $\dot{H}_{2}$ depends on the balance between $z$ and $s(x-x_{R})$; when $z>s(x-x_{R})$, which typically occurs during spiking with $x>x_{R}$, we obtain $\dot{H}_{2}>0$. This mechanism leads to the staircase-like increase of $H_{2}$, representing neuronal adaptation as a slow negative feedback.
@@ 3.3 Lorenz System
Within the framework of NNET, we introduce the following expressions for the Hamiltonians $H_{1}$, $H_{2}$, and the entropy SS [^6] Axenides and Floratos (A&F)[38] study dissipative dynamics by adopting a Helmholtz decomposition of the vector field (rotational part = non-dissipative, gradient part = dissipative), which is formally compatible with the type of splitting used in our discussion. Our NNET framework, however, is formulated as a non-equilibrium thermodynamics: the irreversible contribution is required to take the thermodynamic form $g^{ij}\nabla_{j}S$ with a non-negative dissipation metric, so that the structure of $\dot{S}$ explicitly separates entropy exchange from non-negative entropy production. In contrast, the gradient term in A&F is introduced primarily to represent phase-space volume contraction and is not tied to such thermodynamic constraints. Finally, the reversible/irreversible separation itself is generally non-unique; in NNET this is reflected, for example, in the reparameterization freedom of $H_{1},H_{2}$ (e.g., an $SL(2,\mathbb{R})$-type freedom). More concretely, this redundancy allows one to reshuffle contributions between the Nambu and gradient parts by a compensating transformation of the form $0=\{x^{i},H_{1},Q\}_{\mathrm{NB}}+\partial_{i}(\delta S)$ (for a suitable pair $(Q,\delta S)$), which leaves $\dot{x}^{i}$ unchanged while modifying the split. Accordingly, our representation $(H_{1},H_{2},S)$ is written so as to retain this redundancy.:
@@ 3.4 Chen System
Figure 13 shows the typical attractor of the Lorenz system in three-dimensional phase space. As in the case of the Lorenz system, in Figures 14, 15 we evaluate the intersection time series $H_{1}(t_{k})$ and $S(t_{k})$ on the section $z=z^{\ast}:=2c-a$, corresponding to the components of the zero equilibrium. As $c$ increases, the point sequence transitions from unimodal dispersion to multimodal and broad distribution, suggesting the progression from the destabilization of periodic solutions to the chaotic phase. In particular, the band expansion of $H_{1}$ is consistent with the characteristic driving enhancement of the Chen system (the increase of 2ca2c-a).
@@ 4 Summary and Discussion
We have shown that Nambu Non-equilibrium Thermodynamics (NNET) gives a practical and unified description of far-from-equilibrium dynamics across four paradigmatic systems: the BZ reaction, the Hindmarsh–Rose neuron, and the Lorenz and Chen chaotic flows. In all cases the decomposition into multi-Hamiltonian (incompressible) flow and an entropic (gradient) flow cleanly separates structure-forming dynamics from dissipation.
For the BZ reaction, the dynamics require multiple Hamiltonians together with an explicit entropic term. Along a cycle, entropic and Hamiltonian contributions almost cancel, concentrating entropy changes into short “kicks”; this clarifies why Onsager’s near-equilibrium theory fails here and how oscillations persist despite dissipation. For the Hindmarsh–Rose model, the zero-set conditions for $\dot{H}_{1},\dot{H}_{2},\dot{S}$ identify phase-flip points on the limit cycle. Near these intersections, $H_{1}$ behaves as a pseudo-conserved quantity, which accounts for the timing of spikes, while the slow dynamics of $H_{2}=z$ with $\dot{H}_{2}=r{s(x_{R}-x)+z}$ explain its staircase-like increase.
@@
Although we focused on spatially uniform ODEs, the construction readily extends to spatially extended systems. This suggests new analyses of entropy–Hamiltonian interplay in pattern formation (e.g., Turing[16] and reaction–diffusion[39, 40]) and other oscillatory or chaotic media (e.g., Bray–Liebhafsky (BL) model[41], lasers[42], astrophysics[43], earthquakes[44], engineering applications[45, 46], and psychology[47]). We hope that these results will eventually contribute to a unified understanding of spike transmission in neuroscience or to a new classification principle for chemical reaction systems; these remain important subjects for future work.
@@ Acknowledgments
Appendix Appendix A NNET in summary (for the present 3D applications)
We consider an $n$-dimensional state vector $\bm{x}=(x^{1},\dots,x^{n})$ on a regular region of the state space. Following the axiomatic formulation of NNET[25], the time evolution is decomposed into a reversible part and an irreversible part,
$\dot{x}^{i}=v_{\rm rev}^{i}(\bm{x})+v_{\rm irr}^{i}(\bm{x}).$ (104)
The reversible part is generated by the Nambu bracket with multiple Hamiltonians, while the irreversible part is a gradient flow driven by an entropy-like potential $S$ with a positive definite coefficient (metric) $g^{ij}$.
In the 3D systems studied in this paper ($n=3$), the Nambu bracket is the Jacobian
$\{F,G,H\}:=\frac{\partial(F,G,H)}{\partial(x,y,z)}=\nabla F\cdot(\nabla G\times\nabla H).$ (105)
The NNET evolution law used throughout the main text reads
$\dot{\bm{x}}=\{\bm{x},H_{1},H_{2}\}+\sum_{j=1}^{3}g^{ij}\,\nabla_{j}S,$ (106)
(or its component form with $g^{ij}$). Here “reversible” is used in the thermodynamic-structure sense: the Nambu part is a volume-preserving (incompressible) flow and conserves the Hamiltonians in the absence of dissipation; it is not meant to imply strict time-reversal invariance under all physical constraints (e.g., positivity constraints of variables).
The entropy rate is decomposed as
$\dot{S}=\{S,H_{1},H_{2}\}+\sum_{i,j=1}^{3}(\nabla_{i}S)^{\mathsf{T}}g^{ij}(\nabla_{j}S).$ (107)
The second term is non-negative for positive definite $g^{ij}$ and corresponds to entropy production. The first term is not sign-definite and represents entropy exchange/flux; hence $\dot{S}$ can be negative in open or far-from-equilibrium situations.
The decomposition in Eq. (106) is not unique in general. (i) The Hamiltonians admit reparameterizations $(H_{1},H_{2})\mapsto(\Phi(H_{1},H_{2}),\Psi(H_{1},H_{2}))$. (ii) One may shift contributions between the Nambu and gradient parts unless additional conditions are imposed.
In this work we fix the split by the following practical rule: (1) choose $(H_{1},H_{2})$ so that the Nambu part reproduces the divergence-free component of the given vector field, (2) choose $S$ so that the remaining component is captured by a gradient flow with a positive definite $g^{ij}$, (3) when multiple choices exist, select the simplest polynomial (or lowest-order) $(H_{1},H_{2},S)$ consistent with the model parameters.
Appendix Appendix B Analysis of cycles and spikes
@@
Appendix Appendix C Harmonic oscillators
@@
$\dot{X}^{i}$ $=-\{X^{i},H_{1},H_{2}\}+\frac{\partial S}{\partial X^{i}}.$ (129)
@@
Appendix Appendix D Turning points of spikes

### v4

@@ 3.2 Hindmarsh-Rose model
$\frac{\partial^{2}H_{1}}{\partial x^{2}}+\frac{\partial^{2}H_{1}}{\partial y^{2}}$ $=$ $\color[rgb]{0,0,0}-\color[rgb]{0,0,0}1\color[rgb]{0,0,0}-\color[rgb]{0,0,0}2dx,$ (71)
@@
$\frac{\partial H_{1}}{\partial x}$ $=$ $\color[rgb]{0,0,0}-\color[rgb]{0,0,0}f(x,y),$ (74)
@@
$\frac{\partial f(x,y)}{\partial x}+\frac{\partial g(x,y)}{\partial x}$ $=$ $2dx.$ (76)
@@
The staircase-like increase of $H_{2}$ can be understood directly from
$\dot{H}_{2}=r\{z-s(x-x_{R})\}.$
With $r\ll 1$, so that $z$ acts as a slow variable. The sign of $\dot{H}_{2}$ depends on the balance between $z$ and $s(x-x_{R})$; when $z>s(x-x_{R})$, which typically occurs during spiking with $x>x_{R}$, we obtain $\dot{H}_{2}>0$. This mechanism leads to the staircase-like increase of $H_{2}$, representing neuronal adaptation as a slow negative feedback.
@@ 3.3 Lorenz System
Within the framework of NNET, we introduce the following expressions for the Hamiltonians $H_{1}$, $H_{2}$, and the entropy SS [^7] Axenides and Floratos (A&F)[38] study dissipative dynamics by adopting a Helmholtz decomposition of the vector field (rotational part = non-dissipative, gradient part = dissipative), which is formally compatible with the type of splitting used in our discussion. Our NNET framework, however, is formulated as a non-equilibrium thermodynamics: the dissipative contribution is required to take the thermodynamic form $g^{ij}\nabla_{j}S$ with a non-negative dissipation metric, so that the structure of $\dot{S}$ explicitly separates entropy exchange from non-negative entropy production. In contrast, the gradient term in A&F is introduced primarily to represent phase-space volume contraction and is not tied to such thermodynamic constraints. Finally, the non-dissipative/dissipative separation itself is generally non-unique; in NNET this is reflected, for example, in the reparameterization freedom of $H_{1},H_{2}$ (e.g., an $SL(2,\mathbb{R})$-type freedom). More concretely, this redundancy allows one to reshuffle contributions between the Nambu and gradient parts by a compensating transformation of the form $0=\{x^{i},H_{1},Q\}_{\mathrm{NB}}+\partial_{i}(\delta S)$ (for a suitable pair $(Q,\delta S)$), which leaves $\dot{x}^{i}$ unchanged while modifying the split. Accordingly, our representation $(H_{1},H_{2},S)$ is written so as to retain this redundancy.:
@@ 3.4 Chen System
Figure 13 shows the typical attractor of the Chen system in three-dimensional phase space. As in the Lorenz case, in Figures 14, 15 we evaluate the section values $H_{1}(t_{k})$ and $S(t_{k})$ on the plane $z=z^{\ast}:=2c-a$. As $c$ increases, the distribution changes from a narrow unimodal cluster to broader multimodal bands, indicating the transition from stable motion to chaotic dynamics.
@@ 4 Summary and Discussion
We have shown that Nambu Non-equilibrium Thermodynamics (NNET) gives a practical and unified description of far-from-equilibrium dynamics across four paradigmatic systems: the BZ reaction, the Hindmarsh–Rose neuron, and the Lorenz and Chen chaotic flows. In all cases the decomposition into a multi-Hamiltonian (incompressible) flow and an dissipative gradient flow cleanly separates structure-forming dynamics from dissipation.
For the BZ reaction, the dynamics require multiple Hamiltonians together with an explicit entropic term. Along a cycle, entropic and Hamiltonian contributions almost cancel, concentrating entropy changes into short “kicks”; this clarifies why Onsager’s near-equilibrium theory fails here and how oscillations persist despite dissipation. For the Hindmarsh–Rose model, the zero-set conditions for $\dot{H}_{1},\dot{H}_{2},\dot{S}$ identify phase-flip points on the limit cycle. Near these intersections, $H_{1}$ behaves approximately as a pseudo-conserved quantity, which accounts for the timing of spikes, while the slow dynamics of $H_{2}=z$ with $\dot{H}_{2}=r{s(x_{R}-x)+z}$ explain its staircase-like increase.
@@
Although we focused on spatially uniform ODEs, the construction readily extends to spatially extended systems. This suggests new analyses of entropy–Hamiltonian interplay in pattern formation and other nonlinear media. We hope that the present results will contribute to a more unified understanding of far-from-equilibrium oscillatory, spiking, and chaotic dynamics888e.g., Turing[16], reaction–diffusion[39, 40]) and other oscillatory or chaotic media (e.g., Bray–Liebhafsky (BL) model[41], lasers[42], astrophysics[43], earthquakes[44], engineering applications[45, 46], and psychology[47].
@@ Acknowledgments
Appendix Appendix A Analysis of cycles and spikes
@@
Appendix Appendix B Harmonic oscillators
@@
$\dot{X}^{i}$ $=-\{H_{1},H_{2},X^{i}\}+\frac{\partial S}{\partial X^{i}}.$ (131)
@@
Appendix Appendix C Turning points of spikes